OpenStudy (anonymous):

The hypotenuse of a right triangle is 30 cm and the projection of a leg on it 10.8 cm. Find the other leg.

OpenStudy (anthonyym):

 Use pythagorean theorem

Mehek (mehek14):

Use the Pythagorean Theorem \(a^2+b^2=c^2\) a and b are the legs, and c is the hypotenuse \(10.8^2+b^2=30^2\\30^2-10.8^2=c^2\)

OpenStudy (anonymous):

please help me

Mehek (mehek14):

find \(30^2\) and \(10.8^2\)

jhonyy9 (jhonyy9):

@Mehek14 sorry for my question,but i think that this text wan meaning different what you have understood it there - so the projection of a leg on it 10,8 cm - i think that this mean the projection of a leg on hypotenuse and not the length of a leg ,so than this mean that not can using Pythagorean theorem here - hope i understood it right

jhonyy9 (jhonyy9):

@Kainui your opinion please about ? thank you @Directrix please your opinion ??? thank you

Directrix (directrix):

See attached diagram for visual of projection of leg on hypotenuse.

Directrix (directrix):